OpenStudy (anonymous):

Can anyone simplify? 2{{4(a-4)+16}-{2(4a-3)+1]}

OpenStudy (moongazer):

remember that per form first the operations inside the ( ). then [ ] then { } perform division and multiplication from left to right,and also for addition and multiplication and division first before add and subtraction(MDAS)

OpenStudy (moongazer):

you still there?

OpenStudy (anonymous):

So how did you get so smart????

OpenStudy (anonymous):

=2{{4a-16+16}-{8a-6+1}} =2{(4a)-(8a-5)} =2(4a-8a+5) =2(-4a+5) =-8a+10 you need to use BODMAS rule for this.

OpenStudy (anonymous):

8a +10 was not the right answer

OpenStudy (moongazer):

( ) first. 2{[4(a-4)+16]}-{[2(4a-3)+1]} =2{[4a-16 +16]} - {[8a-6+1]} ={[4a]}-{[8a-5]} =4a-8a-5 =2(-4a-5) =-8a+10 i am so slow.... but i think he is correct

OpenStudy (moongazer):

what is the answeR?

OpenStudy (anonymous):

YEE HAW...Yes you got it

OpenStudy (anonymous):

Damn you are so smart.

OpenStudy (anonymous):

You are NOT 15.

OpenStudy (anonymous):

There is no way someone 15 could be that smart.

OpenStudy (anonymous):

Okay, I only have 4 questions left. And I am relying on you to help me get through this.

OpenStudy (moongazer):

Just study and practice. I am not that smart when i am in grade school and i also had a line of 7 in 1st yr but i realized in the summer break of 3rd year that i need to study and be good because my 2 older sisters are already a graduate and i need to be good and study so I will have a good future so I quit playing computer games and focus in studying it is just the start of our 4th year sorry for my bad grammar i am not that good when it comes to english ^_^ that is where i need help hehehe... just study and have fun

OpenStudy (anonymous):

Okay, this one you will not get and neither will I. Oh, you sweet baby. Thank you so much for doing what you are doing.

OpenStudy (moongazer):

do you have any questions left?

OpenStudy (anonymous):

Yea...it's up there'

OpenStudy (moongazer):

because I am doing already my homework i f you are done ^_^

OpenStudy (anonymous):

Oh, don't let me keep you.

OpenStudy (moongazer):

it is ok it is fun helping you and practicing

OpenStudy (anonymous):

I have 4 questions left